John Smits is married. He and his wife make separate contributions. In this taxable year, John Smits purchased a farm machine for $ 2,000,000 and put it into service. His wife runs her own business and purchased and commissioned a qualified business equipment worth $ 30,000. The “combined dollar limit” is $ 470,000.

As determined by the IRS in its publications 534 and 946, determine the following for each case:

- the amount of depreciation allowed for this taxable year if section 179 is not considered.

- the amount of depreciation allowed for this taxable year if section 179 is considered.
Answer #1

ANSWER:

(a) Depreciation Allowable if 179 not considered: The Depreciation allowable if the sectio 179 is not considered would be equal to the amount of depreciation considered for each asset in a single year, i.e. the Depreciation rate implemented * the cost of the equipment.As Section 179 is not getting applicable here.

(b)Depreciation Allowable if 179 is implemented: If section 179 is applicable here then the amount for maximum depreciation would be 40.15%(This is as per IRS regulation) of the total assets purchased by both John and his wife. So, as the calculation goes 40.15% * (20,00,000+30,000) = 8,15,045, from which the combined dollar limit would be substracted ti get the exact depreciable amount.

Hence, it is = 8,15,045 - 4,70,000 = 3,45,045.

Comment below in case of any doubts.